OpenStudy (anonymous):

Use the substitution method to solve the following system of linear equations: 2x + 2y = 10 x = –y + 5

OpenStudy (anonymous):

A. The solution is (4, –1). B. There are an infinite number of solutions. C. There is no solution. D. The solution is (2, –3).

OpenStudy (anonymous):

x = -y + 5 now sub -y + 5 in for x in the other equation 2x + 2y = 10 2(-y + 5) + 2y = 10 -2y + 10 + 2y = 10 -2y + 2y = 10 - 10 0 = 0 INFINITE SOLUTIONS That is using the substitution method, but if you look at the equations.. x = -y + 5 x + y = 5 2x + 2y = 10 -- divide by 2 x + y = 5 you will see that they are the same line and therefore have infinite solutions.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

sure thing :)